OpenStudy (anonymous):

Solve by the elimination method 5x+2y=-7 7x-3y=17

OpenStudy (anonymous):

let say u want to remove the y so they make them cancel out, i'm shock u still having problem (5x+2y=-7 )(3) (7x-3y=17 )(-2)

OpenStudy (anonymous):

you add them up and you will get x alone

OpenStudy (anonymous):

then simply plug in answer for x for either 1 of the 2 equation and find y

OpenStudy (anonymous):

shouldn't be too hard just "combining equations" a more easier understandable definition for the more technical term elimination method

OpenStudy (anonymous):

isolate one of the variables and then just solve

OpenStudy (anonymous):

sorry is multiply (2) in the 2nd equation not -2

OpenStudy (anonymous):

x= 13/19 is the wrong answer...I got that answer and it was wrong

OpenStudy (anonymous):

why? where did you get that answer?

OpenStudy (anonymous):

I go to online school and when we do our checkpoints it lets us do it until you get the right answer if you want, but before you move onto the next problem it tells you if you got the problem right or not.

OpenStudy (anonymous):

x=13/29?

OpenStudy (anonymous):

14+15 is 29

OpenStudy (anonymous):

If I get the answer wrong 3 times it will give me the right answer but then I will gave to get a new problem and start all over.

OpenStudy (anonymous):

I will give it to you hold on.

OpenStudy (anonymous):

13/29, -134/29

OpenStudy (anonymous):

I got the second part completely wrong...Now I have to start over with a whole new problem

OpenStudy (anonymous):

let me try the second part

OpenStudy (anonymous):

but do u understand that is just removing one of the variables away and solving for the other?

OpenStudy (anonymous):

math is about the solving not the answer

OpenStudy (anonymous):

if x =13/19 5x+2y=-7 5(13/19)+2y=-7 65/19 + 2y =-7 2y=-7-65/19 y=(-133-65)/38 y=-198/38 This is what I get. Now let my check if the solutions you have are right

OpenStudy (anonymous):

x = 13/29 emunrradtvamg you add 14 +15 wrong

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

why dont you just shut up!!

OpenStudy (anonymous):

it was a joke, i did mention it though, but you know it , jesus , she's been posting these question for quite a while

OpenStudy (anonymous):

ok try to help her then. She is having trouble with that. Isnt that obvious?

OpenStudy (anonymous):

Hey the answer you have is right!

OpenStudy (anonymous):

15x+6y=-21 14x-6y=-34 29x=13 x=13/29 to find y 5(13/25)+2y=-7 65/29 + 2y =-7 2y=-7-65/29 y=(-203-65)/58 y=-134/29